OpenStudy (anonymous):

Prove the Trig. Identitiy: (1+tanx)/(1+cotx)=tanx

OpenStudy (anonymous):

1 + sin x ---- cos(x) ------------------------- 1+ cos(x) ------ sin(x) cos(x)+sin(x) sin(x) ------------ * cos(x) ------------- sin(x)+ cos(x) sin(x) -----= tan(x) cos(x)

OpenStudy (anonymous):

How did you get from step one to step two?

OpenStudy (anonymous):

found the common denominator

OpenStudy (anonymous):

division is same as multiplication by reciprocal cos(x)+sin(x) ------------ cos(x) ---------------- -------------- sin(x)+ cos(x) -------------- sin(x) so we flip the denominator cos(x)+sin(x) Sin(x) ------------ * -------------- cos(x) cos(x)+sin(x)

hero (hero):

Still confused?

OpenStudy (anonymous):

Yes, slighltly....?

OpenStudy (anonymous):

Slightly*

OpenStudy (anonymous):

what's seems troubling?

OpenStudy (anonymous):

Where did the 1 in the denominator go from step 1 to step 2?

OpenStudy (anonymous):

common denominator 1 + sin x ---- cos(x) cos(x) + sin(x) ------------- cos(x)

OpenStudy (anonymous):

can i butt in? you know \[\tan(x)=\frac{\sin(x)}{\cos(x)}\] and \[\cot(x)=\frac{\cos(x)}{\sin(x)}\] now replace cosine by "a" sine by "b" and show that \[\frac{1+\frac{b}{a}}{1+\frac{a}{b}}=\frac{b}{a}\]

OpenStudy (anonymous):

cos(x)+sin(x) sin(x) ------------ * cos(x) ------------- sin(x)+ cos(x) This is just throwing me all off. Sorry. :\

OpenStudy (anonymous):

it is pure algebra and has nothing at all to do with sine and cosine

OpenStudy (anonymous):

sin(x)+cos(x)= cos(x)+sin(x) thus they cancel

OpenStudy (anonymous):

so ignore them. it is much easier to write \[\frac{1+\frac{b}{a}}{1+\frac{a}{b}}=\frac{b}{a}\] then the equivalent stuff with a bunch of trig functions. this is my best suggestion for doing these problems. do the algebra with variables and then put the trig functions back in

OpenStudy (anonymous):

i wouldn't even bother with adding this stuff up. i would multiply top and bottom by \[ab\] and clear the fractions.

OpenStudy (anonymous):

\[\frac{1+\frac{b}{a}}{1+\frac{a}{b}}\times \frac{ab}{ab}\] \[\frac{ab+b^2}{ab+a^2}\] \[\frac{b(b+a)}{a(b+a)}=\frac{b}{a}\]

OpenStudy (anonymous):

done in three steps.

OpenStudy (anonymous):

Oh my goodness, @satellite73, thank you so so so very much. (:

OpenStudy (anonymous):

i hope it is clear that this particular problem has nothing whatsoever to do with trigonometry. it would work with numbers, variables, two other functions, anything

OpenStudy (anonymous):

Yes, I was trying way too hard to apply a trig ideas. Inserting a's and b's helped a lot. (: